find the slope and the y-intercept for the following equation of a line 6x+3y=12
Answer by RAY100(1637)

6x +3y =12,,,,,rewrite into slope intercept form,,,y=mx+b, where m=slope, and b= y intercept
.
3y=-6x +12
.
y= -2x + 4,,,,therefore m=(-2) and b=(+4)
.